OpenStudy (anonymous):

6x+y=9.7 4x+2y=9.8 , What is the value of x? please help me :( @-------College Al…

OpenStudy (amistre64):

multiply the top by -2 and add the equations together to start with

OpenStudy (amistre64):

-2(6x+y=9.7) 4x+2y=9.8 -12x-2y=9.7(-2) 4x+2y=9.8 --------------- -8x + 0 = 9.7(-2) + 9.8 x = (9.7(-2) + 9.8)/-8

OpenStudy (amistre64):

once you get rid of the y part, x is at your mercy ....

OpenStudy (anonymous):

ok, do i solve for the x = (9.7(-2) + 9.8)/-8

OpenStudy (amistre64):

you should solve it as you go thru it, but I aint got a calculator handy to do the stuff with. so yes, finish off the math to get an answer :)

OpenStudy (anonymous):

get the value of y from first equation! y=9.7-6x put the value of y in terms of x in the 2nd equation! 4x+2(9.7-6x)=9.8 solving would get the value of x! and further putting the value of x in equ 1 would yeild y

OpenStudy (anonymous):

1.2